# ACDiT: Interpolating Autoregressive Conditional Modeling and Diffusion Transformer

### v1

@@ 5.2 Main Results
Image Generation. We report the FID-50K [26], Inception Score [58], Precision and Recall [35] of ACDiT and baselines in Table 1. Compared with previous autoregressive models and masked generative models utilizing discrete tokens, such as VQGAN, VAR, LlamaGen, and MaskGIT, ACDiT consistently achieves superior performance with lower FID scores at comparable model scales. Notably, ACDiT-XL achieves 2.45 FID scores, outperforming both LlamaGen-XXL and VAR-d20 with similar parameters. Additionally, when compared to the MAR-L variant that does not recompute attention, ACDiT-L significantly improves performance across all metrics. Although MAR-L has lower FID than ACDiT, recomputing attention makes it hard to generalize longer sequence generation. When compared with leading diffusion-based methods, ACDiT also demonstrates competitive performance. For instance, despite not employing full-sequence attention, ACDiT models achieve results close to DiT-XL. In general, these results highlight the distinct advantages of ACDiT over other baselines with the continuous latent representation and KV-Cache. Qualitative results are presented in Fig. 5.
Video Generation. Different from image generation, video inherently includes a temporal dimension, making it more well-suited to autoregressive modeling. The FVD metric on UCF-101 for class-conditional video generation is reported in Table 3. With hybrid AR+Diff architecture, ACDiT-H achieves much lower FVD than other diffusion-based and autoregressive methods, even outperforming MAGVITv2-AR, which utilizes a closed-source, specially designed video tokenizer. In contrast, ACDiT simplifies the process by directly using an open-sourced image VAE. Although MAGVITv2 with masked generative methods has a lower than ACDiT, they rely on ”in-place” operation to generate a video similar to the diffusion model. This constraint limits their ability to generalize to generate longer video generation and build world models. Compared to image generation, ACDiT demonstrates greater potential in modeling long visual sequences. Qualitative results of ACDiT-XL are presented in Fig. 6.
Image Representation. We also assess the capability of ACDiT in image representation, which is essential for building a unified visual understanding and generation model. We finetune ACDiT-XL and DiT-XL on ImageNet using classification loss following the training setting in MAE [24] and report the Top-1 accuracy in Table 4. The accuracy of ACDiT surpasses that of ViT-H, iGPT and DiT-XL. This superior performance over DiT-XL highlights the benefit of incorporating clean latent inputs, which accelerates the model’s ability to learn better representations compared to using only noised latent inputs. Furthermore, ACDiT is on par with MAGE in terms of Top-1 accuracy, while enjoying better generation capabilities than MAGE.
@@ 6 Conclusion
In this paper, we propose ACDiT that interpolates the autoregressive modeling and diffusion transformers. With a simple but novel design of attention mask, ACDiT can achieve autoregressive generation on any length while maintaining a clear latent input potentially for adding a visual understanding task. We demonstrate the performance and efficiency of ACDiT in image and video generation tasks while endowing sufficient generalization by combining the advantages of both autoregressive and diffusion models. We hope ACDiT can shed light on the architectural design of building a unified multimodal model and world model in the future.

### v2

@@ 5.2 Main Results
Image Generation. We report the FID-50K [29], Inception Score [61], Precision and Recall [37] of ACDiT and baselines in Table 2. Compared with previous autoregressive models and masked generative models utilizing discrete tokens, such as VQGAN, VAR, LlamaGen, and MaskGIT, ACDiT consistently achieves superior performance with lower FID scores at comparable model scales. Notably, ACDiT-XL achieves 2.45 FID scores, outperforming both LlamaGen-XXL and VAR-d20 with similar parameters. Additionally, when compared to the MAR-L variant that does not recompute attention, ACDiT-L significantly improves performance across all metrics. Although MAR-L has lower FID than ACDiT, recomputing attention makes it hard to generalize longer sequence generation. Compared with other autoregressive diffusion methods, such as Monoformer [86] and Dart [25], ACDiT has a significantly superior performance. When compared with leading diffusion-based methods, ACDiT also demonstrates competitive performance. For instance, despite not employing full-sequence attention, ACDiT models achieve results close to DiT-XL. In general, these results highlight the distinct advantages of ACDiT over other baselines with the continuous latent representation and KV-Cache. Qualitative results are presented in Appendix.
Video Generation. Different from image generation, video inherently includes a temporal dimension, making it more well-suited to autoregressive modeling. The FVD metric on UCF-101 for class-conditional video generation is reported in Table 3. With hybrid AR+Diff architecture, ACDiT-H achieves much lower FVD than other diffusion-based and autoregressive methods, even outperforming MAGVIT-AR and MAGVITv2-AR, which utilizes a closed-source, specially designed video tokenizer. In contrast, ACDiT simplifies the process by directly using an open-sourced image VAE. Although MAGVIT with masked generative methods has a lower FVD than ACDiT, they rely on ”in-place” operation to generate a video similar to the diffusion model. This constraint limits their ability to generalize to generate longer video generation and build world models. Compared to image generation, ACDiT demonstrates a greater potential in the more complex domain of video generation, where longer visual sequences and temporal information are critical. Qualitative results of ACDiT-XL are presented in Fig. 9.
Image Representation. We assess the capability of ACDiT in image representation, which is essential for building a unified visual understanding and generation model. We finetune ACDiT-XL and DiT-XL on ImageNet using classification loss following the training setting in MAE [27] and report the Top-1 accuracy in Table 4. The accuracy of ACDiT surpasses that of ViT-H, iGPT and DiT-XL. This superior performance over DiT-XL highlights the benefit of incorporating clean latent inputs, which accelerates the model’s ability to learn better representations compared to using only noised latent inputs. Furthermore, ACDiT is on par with MAGE [40] in terms of Top-1 accuracy, while enjoying better generation capabilities.
@@ 6 Conclusion
In this paper, we propose ACDiT that interpolates the autoregressive modeling and diffusion transformers. With a simple but novel design of attention mask, ACDiT can achieve autoregressive generation on any length while maintaining a clear latent input potentially for adding a visual understanding task. By combining the advantages of both autoregressive and diffusion models, we demonstrate the performance and efficiency of ACDiT in both image and video generation tasks. We hope ACDiT can shed light on designing the new visual autoregressive model and building a unified multimodal model in the future.